OpenStudy (anonymous):

a triangle has sides a, b, and c, and the angle opposite the side of length a is obtuse what must be true? a. a^2 + b^2 < c^2 b. a^2 + c^2 a^2

ganeshie8 (ganeshie8):

you may use cosine law : \[a^2 = b^2+c^2-2bc\cos(A)\]

OpenStudy (anonymous):

so how do i find the answer